when CH 3 -CH=CH 2 reacts with Br 2 and then CH 2 I 2 what is the product

when CH3-CH=CH2 reacts with Br2 and then CH2I2 what is the product

Grade:

1 Answers

PASUPULETI GURU MAHESH
8 years ago
CH3-CH=CH2 reacts with Br2 and then CH2I2 AND GIVES SHOWN BELOWTHE GIVEN CH2Br2CH2I2 GIVES CH3-C(I2)-CH3++Br2 CH3-CH=CH2 + CH2Br2CH3-C(I2)-CH3+THERE FORE THE PRODUCT IS